# RadarCam-Depth: Radar-Camera Fusion for Depth Estimation with Learned Metric Scale

### 09 Jan 2024 v1

@@ IV-B Training Details and Evaluation Protocol
When training the RC-Net for quasi-dense scale estimation, we utilize the network architecture of [12] with our transformer module. For the training on nuScenes, with an input image size of $900\times 1600$, the size of the cropped patch during confidence map formation is set to $900\times 288$. For the training on ZJU-4DRadarCam, the input image size is $300\times 1280$, while the patch size is $300\times 100$. We employ the Adam optimizer with $\beta_{1}$=0.9 and $\beta_{2}$=0.999, and a learning rate of $2e^{-4}$ for 50 epochs. Data augmentations, including horizontal flipping, saturation, brightness, and contrast adjustments, are applied with a 0.5 probability. We train our RC-Net for 50 epochs with an NVIDIA RTX 3090 GPU, taking approximately 14 hours with a batch size of 6.
We adopt MiDaS-Small architecture in our SML, where the encoder backbone is initialized with pre-trained ImageNet weights [35], and other layers are randomly initialized. The input data is resized and cropped to $288\times 384$. We use an Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The initial learning rate is set to $2e^{-4}$ and reduced to $5e^{-5}$ after 20 epochs. Training SML for 40 epochs takes about 24 hours with a batch size of 24.
Some widely adopted metrics from the literature are used for evaluating the depth estimations, including mean absolute error (MAE), root mean squared error (RMSE), absolute relative error (AbsRel), squared relative error (SqRel), the errors of inverse depth (iRMSE, iMAE), and $\delta_{1}$ [36].
Figure 7: From left to right are the input images, depth estimaiton of [12], and depth estimation of our RadarCam-Depth on nuScenes benchmark. Ours accurately differentiates the sky region and preserves fine details on object boundaries.
@@ IV-C Evaluation on NuScenes
We evaluate the metric dense depth against $\mathbf{d}_{gt}$ within the range of 50, 70, and 80 meters (see Tab.I). Our proposed RadarCam-Depth outperforms all the compared Radar-Camera methods and surpasses the second best method [12] by a large margin at all ranges. Specifically, we observe 25.6%, 23.4%, and 22.5% reductions in MAE and 20.9%, 20.2%, and 19.4% drops in RMSE for 50m, 70m, and 80m, respectively. Furthermore, our approach obviates the need to aggregate multi-frame Radar point clouds or multi-view images. Our method solely relies on a single-frame image and a Radar point cloud, thus enhancing its real-time capabilities. The ”Radar” and ”Image” columns in Tab.I specify the quantities of Radar point clouds and images used as inputs in various methods.
The images in nuScenes contain large regions corresponding to the sky. For our framework’s monocular depth prediction network, we adopt the sky-sensitive pre-trained model, MiDaS v3.1, which can accurately differentiate the sky from others (see Fig.3). However, since the depth estimations in sky regions are not associated with corresponding LiDAR depth, they are not counted during metric evaluations. Some snapshots of depth estimations from different methods are shown in Fig.7.
@@ IV-D Evaluation on ZJU-4DRadarCam
We follow a similar way to Sec.IV-C for the evaluations on the ZJU-4DRadarCam dataset. For the mono-depth prediction in our framework, we tried both MiDaS v3.1 [3] and DPT-Hybrid [20] models. The evaluations of the metric dense depth estimations from various methods are presented in Tab.II, where our approach at different configurations of mono-depth network and global alignment options are marked in bold. After a comprehensive evaluation, we observe that our methods with the DPT model performs better for depth metrics, while the methods with MiDaS demonstrates higher accuracy for inverse depth metrics. Overall, our proposed methodology exhibits significant improvements compared to existing Radar-Camera methods [12] and [11] (Fig.4(b)). Compared to the second best [12], the best configuration of our method shows 40.2%, 40.1%, and 40.2% reductions in MAE, and 24.3%, 24.6%, and 25.1% drops in RMSE within ranges of 50m, 70m, and 80m, respectively.
@@
TABLE III: Runtime (s)
@@ IV-E1 Transformer Module
We commence our analysis by focusing on the transformer mechanism incorporated within our novel quasi-dense scale estimation module, RC-Net. When the transformer component is disengaged, the architecture is identical to the pre-existing vanilla network, RC-vNet [12]. Our evaluation is conducted on the ZJU-4DRadarCam datasets, and the results are presented in Tab.IV. The comparison results are the error in quasi-dense depth estimation $\mathbf{\hat{d}}_{q}$ against ground truth $\mathbf{d}_{gt}$ within a range of 80 meters. Our RC-Net consistently outperforms RC-vNet across all evaluation metrics, as delineated in Tab.IV. Furthermore, as also indicated in Tab.II, when integrated with the DPT+Var framework, DPT+Var+RC-Net exhibits notable performance superiority over its RC-vNet counterpart.
@@ IV-E2 Global Alignment
Following the discussions in Sec.III-B, we systematically assess the four options for globally aligning the scale of mono-depth predictions. It is worth mentioning that we set a termination criterion of 400 iterations for the GA (RANSAC), at which point we halt the iterative process and select the values of $\hat{s}_{g}$ and $\hat{t}_{g}$ that yield the highest inlier ratio. The evaluation results of $\hat{\mathbf{d}}_{ga}$ are presented in Tab.V within a range of 80 m. Combining the runtime performance in III, the best GA method for ZJU-4D (DPT) is aligning with variable $\hat{s}_{g}$ (Var), and the optimal method for ZJU-4D (MiDaS) is least-squares for $\hat{s}_{g}$ and $\hat{t}_{g}$ (LS). The experiments showcase that estimating $\hat{t}_{g}$ for DPT leads to substantial inverse errors, significantly degrading the performance of the subsequent SML (conducted in inverse space). However, ZJU-4D (MiDaS) require simultaneous estimating $\hat{s}_{g}$ and $\hat{t}_{g}$ to achieve higher accuracy.
@@ V Conclusion
This paper presents a novel method for estimating dense metric depth by integrating monocular depth prediction with the scale from sparse and noisy Radar point cloud. We propose a dedicated four-stage framework that effectively combines the high-fidelity fine details of the image and the absolute scale of Radar data, surmounting the inherent challenges of detail loss and the imprecision of metrics that manifest in existing methods based on the direction fusion of Radar and image data or their encodings. Our experimental findings unequivocally demonstrate a significantly superior performance of the proposed methodology over the compared baseline, as substantiated by both quantitative and qualitative assessments. In general, we introduce a pioneering metric depth estimation solution, which is rigorously validated and suitable for application on fusing cameras with 3D or 4D Radars.

### 19 Mar 2024 v2

@@ IV-B Training Details and Evaluation Protocol
For training the RC-Net on nuScenes, with an input image size of $900\times 1600$, the size of the cropped patch during confidence map formation is set to $900\times 288$. For the training on ZJU-4DRadarCam, the input image size is $300\times 1280$, while the patch size is $300\times 100$. We employ the Adam optimizer with $\beta_{1}$=0.9 and $\beta_{2}$=0.999, and a learning rate of $2e^{-4}$ for 50 epochs. Data augmentations, including horizontal flipping, saturation, brightness, and contrast adjustments, are applied with a 0.5 probability. We train our RC-Net for 50 epochs with an NVIDIA RTX 3090 GPU, taking approximately 14 hours with a batch size of 6.
We adopt MiDaS-Small architecture for our SML, where the encoder backbone is initialized with pre-trained ImageNet weights [35], and other layers are randomly initialized. The input data is resized and cropped to $288\times 384$. We use an Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The initial learning rate is set to $2e^{-4}$ and reduced to $5e^{-5}$ after 20 epochs. Training SML for 40 epochs takes about 24 hours with a batch size of 24.
Some widely adopted metrics from the literature are used for evaluating the depth estimations, including mean absolute error (MAE), root mean squared error (RMSE), absolute relative error (AbsRel), squared relative error (SqRel), the errors of inverse depth (iRMSE, iMAE), and $\delta_{1}$ [36]. To better illustrate our experimental details, the demo video is available at https://youtu.be/JDn0Sua5d9o.
Figure 7: From left to right are the input images, depth estimation of [12], and depth estimation of our RadarCam-Depth on nuScenes benchmark. Ours accurately differentiates the sky region and preserves fine details on object boundaries.
@@ IV-C Evaluation on NuScenes
We evaluate the metric dense depth against $\mathbf{d}_{gt}$ within the range of 50, 70, and 80 meters (see Tab.I). Our proposed RadarCam-Depth outperforms all the compared Radar-Camera methods and surpasses the second best method [12] by a large margin at all ranges. Specifically, we observe 25.6%, 23.4%, and 22.5% reductions in MAE and 20.9%, 20.2%, and 19.4% drops in RMSE for 50m, 70m, and 80m, respectively. We attribute the outstanding performance of RadarCam-Depth to the reasonable monocular prediction $\mathbf{\hat{d}}_{m}$ and our scale learning strategy. Notably, RadarCam-Depth solely relies on a single-frame image and a Radar point cloud, obviating the need to aggregate multi-frame data. The “Radar” and “Image” columns in Tab.I specify the quantities of point clouds and images used as inputs for various methods. For nuScenes dataset, we adopt the sky-sensitive pre-trained model, MiDaS v3.1, as our monocular depth prediction network, which can accurately differentiate the sky from others (see Fig.3). However, since the depth estimations in sky regions are not associated with corresponding LiDAR depth, they are not counted during metric evaluations. Some snapshots of depth estimations from different methods are shown in Fig.7.
@@ IV-D Evaluation on ZJU-4DRadarCam
We follow a similar way to Sec.IV-C for the evaluations on the ZJU-4DRadarCam dataset. For the mono-depth prediction in our framework, we tried both MiDaS v3.1 [3], and DPT-Hybrid [20] models. The evaluations of the metric dense depth estimations from various methods are presented in Tab.II, where our approach at different configurations of mono-depth network and global alignment options are marked in bold. After a comprehensive evaluation, we observe that our methods with the DPT model perform better for depth metrics, while the methods with MiDaS demonstrate higher accuracy for inverse depth metrics. Overall, our proposed methodology exhibits significant improvements compared to existing Radar-Camera methods [12] and [11] (Fig.4(b)). Compared to the second best [12], the best configuration of our method shows 40.2%, 40.1%, and 40.2% reductions in MAE within ranges of 50m, 70m, and 80m, respectively.
@@
TABLE III: Runtime Test of our Modules (s)
@@ IV-E1 Transformer Module
We commence our analysis by focusing on the transformer mechanism incorporated within our novel quasi-dense scale estimation module, RC-Net. When the transformer component is disengaged, the architecture is identical to the pre-existing vanilla network, RC-vNet [12]. Our evaluation is conducted on the ZJU-4DRadarCam dataset, and the results are presented in Tab.IV. The comparison results are the error in quasi-dense depth estimation $\mathbf{\hat{d}}_{q}$ against ground truth $\mathbf{d}_{gt}$ within a range of 80 meters. Our RC-Net consistently outperforms RC-vNet across all evaluation metrics, as delineated in Tab.IV. Furthermore, as indicated in Tab.II, when integrated with the DPT+Var framework, DPT+Var+RC-Net exhibits notable performance superiority over its RC-vNet counterpart.
@@ IV-E2 Global Alignment
Following the discussions in Sec.III-B, we systematically assess the four options for globally aligning the scale of mono-depth predictions. It is worth mentioning that we set a termination criterion of 400 iterations for the GA (RANSAC), at which point we halt the iterative process and select the values of $\hat{s}_{g}$ and $\hat{t}_{g}$ that yield the highest inlier ratio. The evaluation results of $\hat{\mathbf{d}}_{ga}$ are presented in Tab.V within a range of 80 m. Combining the runtime performance in Tab.III, the best GA method for ZJU-4D (DPT) uses variable $\hat{s}_{g}$ (Var), and the optimal method for ZJU-4D (MiDaS) is least-squares for $\hat{s}_{g}$ and $\hat{t}_{g}$ (LS). The experiments showcase that estimating $\hat{t}_{g}$ for DPT leads to substantial inverse errors, significantly degrading the performance of the subsequent SML (conducted in inverse space). However, ZJU-4D (MiDaS) require simultaneous estimating $\hat{s}_{g}$ and $\hat{t}_{g}$ to achieve higher accuracy.
@@ V Conclusion
This paper presents a novel method for estimating dense metric depth by integrating monocular depth prediction with the scale from sparse and noisy Radar point clouds. We propose a dedicated four-stage framework that effectively combines the high-fidelity fine details of the image and the absolute scale of Radar data, surmounting the inherent challenges of detail loss and the imprecision of metrics that manifest in existing methods based on the direction fusion of Radar and image data or their encodings. Our experimental findings unequivocally demonstrate a significantly superior performance of the proposed methodology over the compared baseline, as substantiated by both quantitative and qualitative assessments. In general, we introduce a pioneering metric depth estimation solution, which is rigorously validated and suitable for application on fusing cameras with 3D or 4D Radars. In our future endeavors, we aim to enhance the applicability and effectiveness of our proposed method by leveraging vision foundation models pre-trained with abundant data.